Coffee Company blends coffee type A worth $0.93 per pound with coffee type B worth $1.20 per pound to get 30 pounds of a coffee blen worth $1.02 per pound. How many pounds of each type of coffee does the company use?
Answer by Paul(988)

Let the mixture for $0.93 be x
Let the mixture for $1.2 be y
x+y=30
and y=30-x (subsitution)
0.93x+1.2y=1.02(30)
0.93x+1.2(30-x)=30.6
0.93x-1.2x=30.6-36
-0.27x=-5.4
x=20

y=30-20
y=10
HEnce, the mixture for $0.93 is 20pounds and the mixture for $1.2 is 10 pounds.
Paul.